A borrower agrees to repay a loan of $ 3000 by 15 annual repayments of $500, the first repayment being due after five years. Find the annual yield for this transaction.
1
Expert's answer
2020-10-13T18:57:30-0400

We will assume that the annual rate is "i". Then we have:

"500(\frac{1}{(1+i)^5}+\frac{1}{(1+i)^6}+..+\frac{1}{(1+i)^{20}})=3000" .

we go to step two where we obtain:

"\frac{500}{(1+i)^5}\frac{(1-\frac{1}{(1+i)^{16}})}{1-\frac{1}{(1+i)}}=3000"

From it we receive:


"{(\frac{1}{(1+i)^5}-\frac{1}{(1+i)^{21}})}=6(1-\frac{1}{(1+i)})"


Solving the latter numerically, we obtain:

"\frac{1}{(1+i)}=0.91903"

From the latter we obtain i:

"i\approx0.0881." Which is rounded off to 4 decimal places